OpenStudy (anonymous):

HELP (cot(x)^2)/(csc(x)-1)

OpenStudy (anonymous):

\[(\cot (x)^{2})/(\csc(x)-1)\]

OpenStudy (anonymous):

\[\frac{ \cot(x)^{2} }{ \csc(x)-1 }\] PLEASE SOLVE WILL GIVE POINTS

OpenStudy (anonymous):

I WILL GIVE U A MEDAL I SWEAR

OpenStudy (atlas):

k lets get started

OpenStudy (atlas):

convert the expression into sines and cosines I hope you know cot(x) = cos(x) /sin(x) and cosec(x) = 1/sin(x)

OpenStudy (atlas):

Let me know what you get!

OpenStudy (anonymous):

(cos(x)^2)/(sin(x)^2) so then i did that and changed csc(x)-1 to cot(x) and then I got cos(x)^3/sin(x) in the end

OpenStudy (atlas):

oh no ...........tell again what is the relation ship b/n csc(x) and cot(x)??

OpenStudy (anonymous):

csc(x)-1=cot(x) right?

OpenStudy (anonymous):

Can you tell me what identities I use to solve?

OpenStudy (atlas):

no...........the relation is csc^2(x) - 1 = cot^2(x) :P

OpenStudy (anonymous):

Then what do I do?

OpenStudy (anonymous):

Can you supply the solution?

OpenStudy (atlas):

see u can do the question in a variety of ways.................the best way to do such questions that I found is to convert everything into sine and cos and use a single relation: cos^2x +sin^2x =1

OpenStudy (atlas):

Let us convert everything to sine and cosine

OpenStudy (atlas):

\[ \frac{ \cot^2x }{ \csc x-1 }\] \[\frac{ \cos^2x/\sin^2x }{ 1/sinx -1 }\]

OpenStudy (atlas):

right??

OpenStudy (atlas):

let me drop you a hint after simplfying the above expression you can write cos^2(x) as 1- sin^2(x) which can again be broken down to (1-sin x)(1+sin x)

OpenStudy (anonymous):

I dont know how to simplify further.

OpenStudy (anonymous):

I got (sin(x) - sin(x)^3)(1-sin(x)) divided by sin(x)^2

OpenStudy (anonymous):

\[(\sin(x)-\sin(x)^3-\sin(x)^2+\sin(x)^4)/(\sin(x)^2)\] this is what i got.

OpenStudy (atlas):

Let me show you

OpenStudy (atlas):

Simplifying from where I wrote earlier \[\frac{ \cos^2 x. \sin x }{ \sin^2x(1-sin x) }\]

OpenStudy (atlas):

cancel out the sin x and write cos^2x as (1-sin^2x)

OpenStudy (atlas):

you will get \[\frac{ (1-\sin^2x) }{ \sin x (1-\sin x) }\]

OpenStudy (atlas):

and then i expand (1-sin^2x as) \frac{ (1-sin x)(1+sin x) }{ sin x (1-sin x) }

OpenStudy (atlas):

I hope you can finish off the rest

OpenStudy (anonymous):

Yeah, I really don't get what you did. Anyone care to explain?

OpenStudy (***[isuru]***):

hey buddy... here's an easy method \[\cot ^{2}x + 1 = \csc ^{2}x\] Sooooo \[\cot ^{2}x = \csc ^{2}x -1\] Now substitute that value in to ur expression \[\frac{ \cot ^{2}x}{ csc x - 1 } = \frac{ \csc ^{2} x -1 }{ csc x - 1 }\] Also... u can write \[\csc ^{2}x -1 =( \csc x+1)(\csc x - 1)\] now substitute it again to ur expression.... \[\frac{ \csc ^{2} x -1 }{ \csc x - 1 }\ = \frac{ (\csc x -1)(\csc x +1) }{ (\csc x-1) }\] now cancel out ..... \[(\csc x -1)\] from both numerator and denominator And ur final answer will be...\[\csc x +1\]

OpenStudy (***[isuru]***):

Did ya get it bro ...?

OpenStudy (anonymous):

How did csc(x) become squared?

OpenStudy (anonymous):

OH i get it thank you!

OpenStudy (***[isuru]***):

u r welcome!!